Bbxblue23:

8b -12 + b + 2b + 5 = 180 what is the answer for b?

ArielaQT:

Do you know how to combine like terms?

Bbxblue23:

yep

ArielaQT:

So how would you move -12 and +5 to the right of the equation?

Bbxblue23:

yes

ArielaQT:

It would be the opposite right, so you'd add 12 and subtract 5

ArielaQT:

What's 180 -5 +12?

Bbxblue23:

187

ArielaQT:

So you would be left with, 8b +b + 2b = 187

ArielaQT:

What's 8b +1b +2b?

Bbxblue23:

11b

ArielaQT:

11b = 187.

Bbxblue23:

then i would have to divide them both by 11 right

ArielaQT:

Yep.

ArielaQT:

Giving you b = 17.

Bbxblue23:

so that would make b=17

ArielaQT:

Exactly. Nice work.

Bbxblue23:

thanks bruv

ArielaQT:

No problem.

ArielaQT:

Need anymore help?

IhelpUuHELPme:

sorry it is 17 8b -12 + b + 2b +5 =180 ?b 8b +b+2b = 180+12-5 11b=187 11b 187 b?187/11=17